pokie:

4n ^5 + 3.5n ^5 + (-2.1n^ 5)

pokie:

need help to anwser

Vocaloid:

these are like terms so first step: just add the coefficients 4 + 3.5 + (-2.1) = ?

pokie:

thanks

Vocaloid:

what did you get as your final answer?

pokie:

yes

Vocaloid:

what was your final answer?

pokie:

-16x+24

Vocaloid:

that's not correct again, what is 4 + 3.5 + (-2.1) = ?

pokie:

5.4

Vocaloid:

good, so putting back the variable part gives us 5.4n^5 = your answer

pokie:

thank you